OpenStudy (anonymous):

Express the complex number in trigonometric form. -6 + 6square root of 3i

OpenStudy (jango_in_dtown):

\[-6+6\sqrt{3i}\]

OpenStudy (jango_in_dtown):

??

OpenStudy (anonymous):

Yes but the i isn't under the square root

OpenStudy (jango_in_dtown):

or \[-6+6\sqrt{3}i\]

OpenStudy (jango_in_dtown):

ok \[-6+6\sqrt{3}i=-12(\frac{ 1 }{ 2 }-\frac{ \sqrt{3} }{ 2 }i)\]

OpenStudy (trojanpoem):

x = -6 y = 6sqrt(3) tan theta = 6sqrt(3) / 6 theta = 60 ( in 2nd quarter ) = 120 = 2pi/3 |Z| = sqrt( (-6)^2 + (6sqrt(3))^) = 12 Z = |Z| ( CosTheta + sintheta) = 12 (cos2pi/3 + sin2pi/3)

OpenStudy (jango_in_dtown):

\[=12(\frac{ -1 }{ 2 }+i \sqrt{3}/2)\]

OpenStudy (anonymous):

Thank you @TrojanPoem and @jango_IN_DTOWN!! :)